OpenStudy (anonymous):

If y varies inversely as x and y = 24 when x = 3, what is y when x = 8?

OpenStudy (anonymous):

y=9

OpenStudy (anonymous):

\[\frac{24}{y}=\frac{8}{3}\]24*3=8y 72/8=y 9=y